Determine (OH), ph and pOH of a solution [H]+ with 2.5 X 10 -8 M


1
Expert's answer
2021-03-22T06:31:23-0400

H+ ion Concentration = 2.5×10-8 + 10 -7



So , H+ ion Concentration = 1.25 × 10-7

So, PH = -log 1.25 × 10-7 = 6.9( approx ).. ans.


POH = 14 - 6.9 = 7.1 ( approx ) .. ans.


Concentration of OH- = 10-7.1 ( approx ) ans.